For each x ∈ R, let us denote by C(x) the least integer greater than or equal to x.
For example, C(1) = 1, C(−√2) = −1. In other words, C(x) is the unique integer
satisfying C(x) − 1 < x ≤ C(x).
(1) Draw the graph of the function C(x) for x ∈ [−2, 2].
(2) Prove that C(x) is continuous at all non-integer points of R.
(3) Prove that C(x) is discontinuous at all integer points of R.
1
Expert's answer
2019-04-03T13:00:53-0400

Let x be a real number. The ceiling function of x, denoted by "\lceil x\rceil", is the smallest integer that is larger than or equal to x.

For "x\in[-2, 2]"


"x=-2, f(x)=\lceil x\rceil=-2"

"-2\lt x \le-1, f(x)=\lceil x\rceil=-1"

"-1\lt x \le0, f(x)=\lceil x\rceil=0"

"0\lt x \le1, f(x)=\lceil x\rceil=1"

"1\lt x \le2, f(x)=\lceil x\rceil=2"




For "a_i\in(i, i+1), i\in \Z"


"\lim\limits_{x\to a_i}f(x)=i+1, f(a)=i+1, x\in(i, i+1)"

Therefore, the function "f(x)" is continuous at "x\in \R, x\notin \Z"



For "a_i= i\in \Z:"

for "x\in(i-1, i)"


"\lim\limits_{x\to a_i}f(x)=i,"

for "x\in(i, i+1)"


"\lim\limits_{x\to a_i}f(x)=i+1."

Then


"\lim\limits_{x\to a_i}f(x)=does \ not \ exist"

Therefore, the function "f(x)" is discontinuous at "x\in \Z."